# MapAnything: Universal Feed-Forward Metric 3D Reconstruction map-anything.github.io

### v2

@@ Abstract
We introduce MapAnything, a unified transformer-based feed-forward model that ingests one or more images along with optional geometric inputs such as camera intrinsics, poses, depth, or partial reconstructions, and then directly regresses the metric 3D scene geometry and cameras. MapAnything leverages a factored representation of multi-view scene geometry, i.e., a collection of depth maps, local ray maps, camera poses, and a metric scale factor that effectively upgrades local reconstructions into a globally consistent metric frame. Standardizing the supervision and training across diverse datasets, along with flexible input augmentation, enables MapAnything to address a broad range of 3D vision tasks in a single feed-forward pass, including uncalibrated structure-from-motion, calibrated multi-view stereo, monocular depth estimation, camera localization, depth completion, and more. We provide extensive experimental analyses and model ablations demonstrating that MapAnything outperforms or matches specialist feed-forward models while offering more efficient joint training behavior, thus paving the way toward a universal 3D reconstruction backbone.
@@ 1 Introduction
While prior feed-forward work has approached the different tasks disjointly or by not leveraging all the available input modalities, we present a unified end-to-end model for diverse 3D reconstruction tasks. Our method MapAnything can be used to solve the most general uncalibrated SfM problem as well as various combinations of sub-problems, like calibrated SfM or multi-view stereo, monocular depth estimation, camera localization, metric depth completion, etc. To enable the training of such a unified model, we: (1) introduce a flexible input scheme that supports various geometric modalities when available, (2) propose a suitable output space that supports all of these diverse tasks, and (3) discuss flexible dataset aggregation and standardization.
MapAnything’s key insight to address these challenges is the use of a factored representation of multi-view scene geometry. Instead of directly representing the scene as a collection of pointmaps, we represent the scene as a collection of depthmaps, local raymaps, camera poses, and a metric scale factor that upgrade local reconstructions into a globally consistent metric frame. We use such a factored representation to represent both the outputs and (optional) inputs for MapAnything, allowing it to take advantage of auxiliary geometric inputs when available. For example, robotic applications [20, 28, 16, 1] may have knowledge of camera intrinsics (rays) and/or extrinsics (pose). Finally, a significant benefit of our factored representation is that it allows for MapAnything to be effectively trained from diverse datasets with partial annotations, for example, datasets that may be annotated with only non-metric “up-to-scale” geometry. In summary, we make the following main contributions:
@@ 2 Related Work
In contrast to the traditional approach of designing expert methods for distinct reconstruction tasks, recent efforts have shown great promise in solving them jointly with a single feed-forward architecture. Early works like DeMoN [60], DeepTAM [88] or DeepV2D [56] explored this direction with CNNs but did not match the performance of classical expert models. Enabled by advances in deep learning, recent methods like PF-LRM [67], RayDiffusion [85], DUSt3R [72], VGGSfM [65] or VGGT [66] scale up transformers on large amounts of data. Despite this breakthrough, these methods are still limited to a subset of 3D reconstruction tasks with fixed inputs and output modalities, a small or fixed number of views, or they only work well in relatively constrained, typically object-centric, scenarios. With MapAnything, we overcome these limitations by designing a geometrically grounded and flexible architecture that supports heterogeneous input and output modalities for any number of input views.
Figure 2: Overview of the MapAnything Architecture. Given NN visual and optional geometric inputs, the model first encodes the images and the factored representation of the geometric inputs into a common latent space where the patch features (for images, rays & depth) and broadcasted global features (for translation, rotation, pose scale across all pose inputs & depth scale local to each frame) are summed together. Then, a fixed reference view embedding is added to the first view’s features and a single learnable scale token is appended to the set of $N$ view patch tokens. These tokens are then input into an alternating-attention transformer. We use a single DPT to decode the NN view patch tokens into $N$ dense outputs local to all the views. A single average pooling-based pose head also uses the NN view patch tokens to predict NN poses in the frame of view 1. Lastly, while these predictions exist in an up-to-scale space, the model passes the scale token through an MLP to predict the metric scaling factor, which when coupled with the other predictions, provides the dense metric 3D reconstruction.
@@
DUSt3R and its metric follow-up MASt3R [30] predict a coupled scene representation (i.e., cameras, poses, and geometry are parameterized by a pointmap and need to be recovered post-hoc) and need expensive post-processing & symmetric inference to perform multi-view unconstrained SfM. Follow-up work [10, 39, 11, 43] integrates MASt3R’s outputs into classical SfM and SLAM pipelines in a more principled manner. Recent works like Spann3R [64], CUT3R [68], and MUSt3R [6] remove the need for classical optimization and enable multi-view reconstruction via latent state memory in transformers. However, these works do not yet match the performance of traditional optimization on top of predicted two-view outputs from MASt3R [10, 39].
Recently, MV-DUSt3R+ [55] and VGGT [66] demonstrate multi-view inference by extending the DUSt3R architecture for multi-view reconstruction. Likewise, Reloc3r [9] focuses on camera re-localization and directly predicts multi-view camera poses. MV-DUSt3R+ achieves this by parallelizing the cross-attention transformer to support different reference views, leading to a significant increase in compute, while VGGT employs an alternating attention transformer to predict multi-view pointmaps, depth, pose, and features for tracking. FASt3R [78] uses positional encoding for long-sequence inference in LLMs for global attention trained on a few views to work on a larger number of views. More recently, $\pi^{3}$ [74] finetunes VGGT to remove the use of the first input frame as reference coordinate.
Across both MV-DUSt3R+ and FASt3R, the prediction is a coupled scene representation and cannot deal with heterogeneous inputs. As shown in FASt3R, for the multi-view setup, the dense geometry prediction capabilities of the model are impacted by the pose estimation across non-visible views (see their Table 5 & Section 5.1). To alleviate this issue, FASt3R predicts redundant pointmaps across all views with a dedicated DPT head for global and local pointmap prediction. Likewise, VGGT also predicts multiple redundant quantities through two separate branches, one for point maps and one for cameras and depth. While concurrent work, $\pi^{3}$ [74], finetunes VGGT to remove this redundancy by predicting up-to-scale decoupled local pointmaps and global pose, we find this design choice to be sub-optimal (see Table˜5(a)). In contrast, MapAnything directly predicts a completely factored representation, i.e., local ray directions, depth along the ray, global camera pose for all views, and a single metric scaling factor for the scene. In this formulation, the task of predicting ray directions (akin to camera calibration) and depth-along-ray estimation are per-view and thus can be predicted from a single dense prediction head.
@@
While not explicitly used for feed-forward 3D reconstruction, like in MapAnything, quantities such as ray directions, origins, and depth maps have been explored as conditioning inputs for tasks like novel-view synthesis [26, 36, 75, 89], or diffusion-based image generation [38, 86]. Similarly, Align3R [35] uses monocular depth priors as conditioning for dynamic video depth estimation. Taskonomy [84] explored the benefits of multi-task learning for improved vision task performance. Later works like MultiMAE [3] build on these insights and devise auto-encoders to support flexible combination of heterogeneous inputs; however, this is not suitable for solving 3D reconstruction tasks. Pow3R [25] was the first to incorporate known priors as inputs to feed-forward 3D reconstruction. In contrast to us, Pow3R only supports two pinhole camera images with a single focal length and centered principal point. Furthermore, Pow3R builds on top of DUSt3R and cannot condition on metric scale information. In contrast, MapAnything supports any number of input views and has a flexible input parameterization to support metric scale and any camera with a central projection model.
@@ 3 MapAnything
where $m\!\in\!\mathbb{R}$ is the predicted global metric scaling factor, and for each view $i$, $R_{i}\!\in\!\mathbb{R}^{3\times H\times W}$ are the predicted local ray directions, $\tilde{D}_{i}\!\in\!\mathbb{R}^{1\times H\times W}$ are the ray depths in a up-to-scale space (indicated by the tilde), and $\tilde{P}_{i}\!\in\!\mathbb{R}^{4\times 4}$ is the pose of image $\hat{I}_{i}$ in the frame of image $\hat{I}_{1}$, represented as quaternion $Q_{i}\!\in\!\mathrm{SU}(2)$ and up-to-scale translation $\tilde{T}_{i}\!\in\!\mathbb{R}^{3}$. We can further use this factored output to get the up-to-scale local point maps (3D points corresponding to each pixel) as $\tilde{L}_{i}\!=\!R_{i}\!\cdot\!\tilde{D}_{i}\!\in\!\mathbb{R}^{3\times H\times W}$. Then, leveraging the rotation matrix ${O}_{i}\!\in\!\mathrm{SO}(3)$ (obtained from $Q_{i}$) and up-to-scale translation, we can compute the $N$-view up-to-scale point maps in world frame as $\tilde{X}_{i}\!=\!{O}_{i}\!\cdot\!\tilde{L}_{i}\!+\!T_{i}$. The final metric 3D reconstruction for the $N$ input views (in the frame of image 1) is given by ${X}^{\text{metric}}_{i}\!=\!m\!\cdot\!\tilde{X}_{i}$ for $i\!\in\![1,N]$.
Figure 3: Auxiliary geometric inputs improve feed-forward performance of MapAnything. (Top) While MapAnything & other baselines using 100 input images show duplication of 3D structure, when provided with the camera calibration and poses, the 3D reconstruction significantly improves, showcasing aligned geometry. (Middle) MapAnything using images only as input showcases non-precise metric scale estimation on ETH3D (a zero-shot dataset). However, when the calibration and metric poses are provided as additional input, the estimated metric scale significantly improves and approximately matches the ground truth. (Bottom-Left) We showcase that MapAnything is able to leverage a sparse metric pointcloud as input to perform dense metric depth completion. (Bottom-Right) Despite not being trained for object-centric data, we showcase how the scene geometry and cameras change based on the input provided.
@@ 3.1 Encoding Images & Geometric Inputs
Given the $N$ visual inputs and optional dense geometric inputs, we first encode them into a common latent space. For images, we use DINOv2 (Apache 2.0) [41]. Amongst a wide variety of pre-trained options, such as CroCov2 [76], DUSt3R’s image encoder [72], RADIO [47, 15], and random-init linear patchification, we find DINOv2 to be optimal in terms of downstream performance, convergence speed, and generalization (especially when finetuned with a small learning rate). We use the final layer normalized patch features from DINOv2 ViT-L, $F_{\text{I}}\!\in\!\mathbb{R}^{1024\times H/14\times W/14}$.
@@
MapAnything can also encode other geometric quantities. Before feeding these geometric quantities to our network, we factorize them to enable training and inference across both metric and up-to-scale quantities. In particular, when provided, the ray depths are first decoupled into average per-view depth $\hat{z}_{di}\!\in\!\mathbb{R}^{+}$ and normalized ray depths $\hat{D}_{i}/\hat{z}_{di}$. Furthermore, when translations $\hat{\mathcal{T}}$ are provided, MapAnything computes the pose scale as the average distance to the world frame, $\hat{z}_{\text{p}}\!=\!\frac{1}{\left|S_{t}\right|}\sum_{i\in{S_{t}}}\lVert\hat{T}_{i}\rVert$. This pose scale is used as the same input for all frames with input translation and is also used to get the normalized translations $\hat{T}_{i}/\hat{z}_{\text{p}}$. Since we are interested in effectively exploiting the metric scale information from geometric inputs (Figure˜3), MapAnything only uses the pose scale and depth scales when the poses and depths provided for specific frames are metric. Since the metric scale values can be large and drastically vary across scene sizes, we log-transform scales before encoding them.
@@
We encode ray directions and normalized ray depths using a shallow convolutional encoder [38], where the spatial resizing only happens once with a pixel unshuffle of size 14. This projects the dense geometric inputs into the same spatial and latent dimension as the DINOv2 features, i.e., $F_{\text{R}},F_{\text{D}}\!\in\!\mathbb{R}^{1024\times H/14\times W/14}$. For the global non-pixel quantities, i.e, rotations (represented as unit quaternions), translation directions, depth and pose scales, we use a 4-layer MLP with GeLU activations to project the quantities to features $F_{\text{Q}},F_{\text{T}},F_{\hat{z}_{d}},F_{\hat{z}_{\text{p}}}\!\in\!\mathbb{R}^{1024}$. Once all input quantities are encoded, they are passed through layer normalization, summed together, and followed by another layer normalization to obtain the final per-view encodings for each input view. These are then flattened into tokens $F_{\text{E}}\!\in\!\mathbb{R}^{1024\times({HW}/{256})}$.
@@
We append a single learnable scale token to the set of $N$ view patch tokens and input the tokens into a multi-view transformer to allow the information across multiple views to attend to each other and propagate information. We use a 24 layer alternating-attention transformer [66] with 12 blocks of multi-headed attention (12 heads), a latent dimension of 768 and an MLP ratio of 4, which is randomly initialized. To distinguish the reference view (i.e., the first one), we add a constant reference view embedding to the set of patch tokens corresponding to view 1. For further simplicity, we do not use any Rotary Positional Embedding (RoPE) [54]. We find that the patch-level positional encoding from DINOv2 suffices, and RoPE tends to lead to unnecessary biases, given that it was originally applied in every attention layer.

### v3

@@ Abstract
We introduce MapAnything, a unified transformer-based feed-forward model that ingests one or more images along with optional geometric inputs such as camera intrinsics, poses, depth, or partial reconstructions, and directly regresses the metric 3D scene geometry and cameras. MapAnything leverages a factored representation of multi-view scene geometry, i.e., a collection of depth maps, local raymaps, camera poses, and a metric scale factor that effectively upgrades local reconstructions into a globally consistent metric frame. Standardizing the supervision and training across diverse datasets, along with flexible input augmentation, enables MapAnything to address a broad range of 3D vision tasks in a single feed-forward pass, including uncalibrated structure-from-motion, calibrated multi-view stereo, monocular depth estimation, camera localization, depth completion, and more. We provide extensive experimental analyses and model ablations demonstrating that MapAnything outperforms or matches specialist feed-forward models while offering more efficient joint training behavior, thus paving the way toward a universal 3D reconstruction backbone.
@@ 1 Introduction
While prior feed-forward work has approached the different tasks separately or by not leveraging all the available input modalities, we present a unified end-to-end model for diverse 3D reconstruction tasks. Our method MapAnything can be used to solve the most general uncalibrated SfM problem as well as various combinations of sub-problems, such as calibrated SfM or multi-view stereo, monocular depth estimation, camera localization, and metric depth completion. To enable the training of such a unified model, we: (1) introduce a flexible input scheme that supports various geometric modalities when available, (2) propose a suitable output space that supports all of these diverse tasks, and (3) discuss flexible dataset aggregation and standardization.
MapAnything’s key insight to address these challenges is the use of a factored representation of multi-view scene geometry. Instead of directly representing the scene as a collection of pointmaps, we represent the scene as a collection of depth maps, local raymaps, camera poses, and a metric scale factor that upgrade local reconstructions into a globally consistent metric frame. We use such a factored representation to represent both the outputs and (optional) inputs for MapAnything, allowing it to take advantage of auxiliary geometric inputs when available. For example, robotic applications [19, 27, 15, 1] may have knowledge of camera intrinsics (rays) and/or extrinsics (poses). Finally, a significant benefit of our factored representation is that it allows MapAnything to be effectively trained from diverse datasets with partial annotations, for example, datasets that may be annotated with only non-metric “up-to-scale” geometry. In summary, we make the following main contributions:
@@ 2 Related Work
In contrast to the traditional approach of designing specialized methods for distinct reconstruction tasks, recent efforts have shown great promise in solving them jointly with a single feed-forward architecture. Early works like DeMoN [61], DeepTAM [87] or DeepV2D [57] explored this direction with CNNs but did not match the performance of classical expert models. Enabled by advances in deep learning, recent methods like PF-LRM [68], RayDiffusion [84], DUSt3R [72], VGGSfM [66], and VGGT [67] scale up transformers on large amounts of data. Despite this breakthrough, these methods are still limited to a subset of 3D reconstruction tasks with fixed inputs and output modalities, a small or fixed number of views, or they only work well in relatively constrained, typically object-centric, scenarios. With MapAnything, we overcome these limitations by designing a geometrically grounded and flexible architecture that supports heterogeneous input and output modalities for any number of input views.
Figure 2: Overview of the MapAnything Architecture. Given $N$ visual and optional geometric inputs, the model first encodes the images and the factored representation of the geometric inputs into a common latent space where the patch features (for images, rays & depth) and broadcasted global features (for translation, rotation, pose scale across all pose inputs & depth scale local to each frame) are summed together. Then, a fixed reference view embedding is added to the first view’s features and a single learnable scale token is appended to the set of $N$ view patch tokens. These tokens are then input into an alternating-attention transformer. We use a single DPT to decode the NN view patch tokens into $N$ dense outputs local to all the views. A single average pooling-based pose head also uses the NN view patch tokens to predict NN poses in the frame of view 1. Lastly, while these predictions exist in an up-to-scale space, the model passes the scale token through an MLP to predict the metric scaling factor, which when coupled with the other predictions, provides the dense metric 3D reconstruction.
@@
DUSt3R and its metric follow-up MASt3R [29] predict a coupled scene representation (i.e., cameras, poses, and geometry are parameterized by a pointmap and need to be recovered post hoc) and require expensive post-processing & symmetric inference to perform multi-view unconstrained SfM. Follow-up work [10, 39, 11, 44] integrates MASt3R outputs into classical SfM and SLAM pipelines in a more principled manner. Recent works like Spann3R [65], CUT3R [69], and MUSt3R [6] remove the need for classical optimization and enable multi-view reconstruction via latent state memory in transformers. However, these works do not yet match the performance of traditional optimization applied to predicted two-view outputs from MASt3R [10, 39].
Recently, MV-DUSt3R+ [56] and VGGT [67] demonstrate multi-view inference by extending the DUSt3R architecture for multi-view reconstruction. Likewise, Reloc3r [8] focuses on camera re-localization and directly predicts multi-view camera poses. MV-DUSt3R+ achieves this by parallelizing the cross-attention transformer to support different reference views, leading to a significant increase in computation, while VGGT employs an alternating attention transformer to predict multi-view pointmaps, depth, pose, and features for tracking. FASt3R [78] uses positional encoding for long-sequence inference in LLMs for global attention trained on a few views to work on a larger number of views. More recently, $\pi^{3}$ [74] fine-tunes VGGT to remove the use of the first input frame as reference coordinate.
In both MV-DUSt3R+ and FASt3R, the prediction is a coupled scene representation and cannot handle heterogeneous inputs. As shown in FASt3R, for the multi-view setup, the dense geometry prediction capabilities of the model are impacted by the pose estimation across non-visible views (see their Table 5 & Section 5.1). To alleviate this issue, FASt3R predicts redundant pointmaps across all views with a dedicated DPT head for global and local pointmap prediction. Likewise, VGGT also predicts multiple redundant quantities through two separate branches, one for pointmaps and one for cameras and depth. While concurrent work, $\pi^{3}$ [74], fine-tunes VGGT to remove this redundancy by predicting up-to-scale decoupled local pointmaps and global pose, we find this design choice to be sub-optimal (see Table˜5(a)). In contrast, MapAnything directly predicts a completely factored representation, i.e., local ray directions, depth along the ray, global camera pose for all views, and a single metric scaling factor for the scene. In this formulation, the task of predicting ray directions (akin to camera calibration) and depth-along-ray estimation are per-view and thus can be predicted from a single dense prediction head.
@@
While not explicitly used for feed-forward 3D reconstruction like in MapAnything, quantities such as ray directions, origins, and depth maps have been explored as conditioning inputs for tasks like novel-view synthesis [24, 35, 75, 88], diffusion-based image generation [38, 85], dynamic video depth estimation [34], or 3D object shape completion [9]. Taskonomy [83] explored the benefits of multi-task learning for improved vision task performance. Later works like MultiMAE [3] build on these insights and devise auto-encoders to support flexible combination of heterogeneous inputs; however, this is not suitable for solving 3D reconstruction tasks. Pow3R [23] was the first work to incorporate known priors as inputs to feed-forward 3D reconstruction. In contrast to us, Pow3R only supports two pinhole camera images with a single focal length and centered principal point. Furthermore, Pow3R builds on top of DUSt3R and cannot condition on metric scale information. In contrast, MapAnything supports any number of input views and has a flexible input parameterization that supports metric scale and any camera with a central projection model.
@@ 3 MapAnything
where $m\!\in\!\mathbb{R}$ is the predicted global metric scaling factor, and for each view $i$, $R_{i}\!\in\!\mathbb{R}^{3\times H\times W}$ are the predicted local ray directions, $\tilde{D}_{i}\!\in\!\mathbb{R}^{1\times H\times W}$ are the ray depths in a up-to-scale space (indicated by the tilde), and $\tilde{P}_{i}\!\in\!\mathbb{R}^{4\times 4}$ is the pose of image $\hat{I}_{i}$ in the frame of image $\hat{I}_{1}$, represented as quaternion $Q_{i}$ and up-to-scale translation $\tilde{T}_{i}\!\in\!\mathbb{R}^{3}$. We can further use this factored output to get the up-to-scale local pointmaps (3D points corresponding to each pixel) as $\tilde{L}_{i}\!=\!R_{i}\!\cdot\!\tilde{D}_{i}\!\in\!\mathbb{R}^{3\times H\times W}$. Then, leveraging the rotation matrix ${O}_{i}$ (obtained from $Q_{i}$) and up-to-scale translation, we can compute the $N$-view up-to-scale pointmaps in world frame as $\tilde{X}_{i}\!=\!{O}_{i}\!\cdot\!\tilde{L}_{i}\!+\!\tilde{T}_{i}$. The final metric 3D reconstruction for the $N$ input views (in the frame of image $I_{1}$) is given by ${X}^{\text{metric}}_{i}\!=\!m\!\cdot\!\tilde{X}_{i}$ for $i\!\in\![1,N]$.
@@ 3.1 Encoding Images & Geometric Inputs
Given $N$ visual inputs and optional dense geometric inputs, we first encode them into a common latent space. For images, we use DINOv2 (Apache 2.0) [42]. Among a wide variety of pre-trained options, such as CroCov2 [76], DUSt3R’s image encoder [72], RADIO [48, 14], and random-init linear patchification, we find DINOv2 to be optimal in terms of downstream performance, convergence speed, and generalization (especially when fine-tuned with a small learning rate). We use the $24$th layer normalized patch features from DINOv2 ViT-G, $F_{\text{I}}\!\in\!\mathbb{R}^{1536\times H/14\times W/14}$.
MapAnything can also encode other geometric quantities. Before feeding these geometric quantities to our network, we factorize them to enable training and inference across both metric and up-to-scale quantities. To support use cases where only rotation or translation might be individually present (for e.g., IMU & GPS priors) and to deal with the entanglement of translation with scale, we encode rotation and translation separately. Furthermore, since we don’t assume depth & pose to always be provided together as input, we decouple their normalization (note that this is separate from the training objective where we normalize predicted depth & pose together since we want multi-view consistency).
@@
In particular, when provided, the ray depths are first decoupled into average per-view depth $\hat{z}_{di}\!\in\!\mathbb{R}^{+}$ and normalized ray depths $\hat{D}_{i}/\hat{z}_{di}$. Furthermore, when translations $\hat{\mathcal{T}}$ are provided, MapAnything computes the pose scale as the average distance to the world frame, $\hat{z}_{\text{p}}\!=\!\frac{1}{\left|S_{t}\right|}\sum_{i\in{S_{t}}}\lVert\hat{T}_{i}\rVert$. This pose scale is used as the same input for all frames with input translation and is also used to get the normalized translations $\hat{T}_{i}/\hat{z}_{\text{p}}$. Since we are interested in effectively exploiting the metric scale information from geometric inputs, MapAnything only uses the pose scale and depth scales when the poses and depths provided for specific frames are metric. Furthermore, the metric scale values can be large and drastically vary across scene sizes, hence, we log-transform scales before encoding them.
@@
We encode ray directions and normalized ray depths using a shallow convolutional encoder [38], where the spatial resizing only happens once with a pixel unshuffle of size 14. This projects the dense geometric inputs into the same spatial and latent dimension as the DINOv2 features, i.e., $F_{\text{R}},F_{\text{D}}\!\in\!\mathbb{R}^{1536\times H/14\times W/14}$. For the global non-pixel quantities, i.e, rotations (represented as unit quaternions), translation directions, depth and pose scales, we use a 4-layer MLP with GeLU activations to project the quantities to features $F_{\text{Q}},F_{\text{T}},F_{\hat{z}_{d}},F_{\hat{z}_{\text{p}}}\!\in\!\mathbb{R}^{1536}$. Once all input quantities are encoded, they are passed through layer normalization, summed together, and followed by another layer normalization to obtain the final per-view encodings for each input view. These are then flattened into tokens $F_{\text{E}}\!\in\!\mathbb{R}^{1536\times({HW}/{256})}$.
@@
We append a single learnable scale token to the set of $N$ view patch tokens and input the tokens into a multi-view transformer to allow information across multiple views to attend to each other and propagate. We use a 16-layer alternating-attention transformer [67] with 24 heads of multi-headed attention, a latent dimension of 1536 and an MLP ratio of 4, initialized using the last 16 layers of DINOv2 ViT-G [41, 31]. To distinguish the reference view (i.e., the first one), we add a constant reference view embedding to the set of patch tokens corresponding to view $I_{1}$. For simplicity, we do not use Rotary Positional Embedding (RoPE) [55]. We find that the patch-level positional encoding from DINOv2 suffices, and RoPE leads to unnecessary biases, given that it was originally applied in every attention layer.